OpenStudy (anonymous):

Using the factors (-1) and (x-4) find the remaining factor(s) of f(x) = x^3 - 2x^2 - 11x + 12 and write the polynomial as a product of linear factors

OpenStudy (mathmale):

One way to approach this problem would be to divide the given polynomial, f(x), by x-1 and x-4 (separately). Doing this will leave you with the one remaining factor. We know that there are 3 factors because the order of the poly. is 3. I'd suggest using synth. div. if you know it. Try 4 | 1 -2 -11 12 --------------- 1

OpenStudy (mathmale):

Otherwise, try dividing (x-4) into x^3 -2x^2 - 11x +12.

OpenStudy (anonymous):

I come up with (x*3) as my other factor so the product if linear factors would be. (-1)(x+3)(x-4)

OpenStudy (whpalmer4):

are you sure it isn't supposed to be \((x-1)\) as a factor, not just \(-1\)?

OpenStudy (anonymous):

Yes it is (x-1)

OpenStudy (mathmale):

Then it appears that you have completely answered this question. Your polynomial can now be written as f(x) = x^3 - 2x^2 - 11x + 12 = (x-1)(x+3)(x-4). Nice work!